OpenStudy (anonymous):

Question for tosin! What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (anonymous):

@tosin

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

how do i give more medals?

OpenStudy (anonymous):

@tosin

OpenStudy (anonymous):

–4q + 8 = –2q – 16

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

–4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (anonymous):

q=12... D

OpenStudy (anonymous):

Thanks, man, your the best!

OpenStudy (anonymous):

i got you fam.

OpenStudy (anonymous):

how do i give more medals xD

OpenStudy (anonymous):

it would have to be a separate question to award another medal

OpenStudy (anonymous):

oh danger

OpenStudy (anonymous):

danget*

OpenStudy (anonymous):

its cool , lol